# On geometric spectral functionals

### v1

@@ 4 Chiral spectral functionals
$\begin{split}\mathscr{W}\!\mathrm{res}(\chi D^{-n+2})-\mathscr{W}\!\mathrm{res}(\chi D_{0}^{-n+2})&=\frac{\nu_{n-1}}{24}(n-2)\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\,\bigl{[}\chi(-12B_{0}^{2}-12i\gamma^{a}B_{a}\\ &\hskip 85.35826pt+6i\{\gamma^{a},B_{a}\}+3\{\gamma^{a},B\}\{\gamma^{a},B\})\bigr{]}\\ &=\frac{\nu_{n-1}}{8}(n-2)\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\,\bigl{[}\chi(-4B_{0}^{2}-4i\gamma^{a}B_{a}+\{\gamma^{a},B\}\{\gamma^{a},B\})\bigr{]}\\ &=-\frac{i}{2}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(\chi\gamma^{a}B_{a}).\end{split}$ (4.13)
@@ 4.1 The spin Dirac operator
$\mathscr{g}_{D_{0}}^{\chi}(u,w)=\begin{cases}4\pi i{2^{m}}\int_{M}\mathrm{dvol}_{g}u_{a}w_{b}\varepsilon^{ab},\quad n=2,\\ 0,\quad n>2.\end{cases}$ (4.14)
Proof 4.9.
It follows from Proposition 2.1.
@@
$\mathscr{T}_{D}^{\chi}(u,v,w)=2^{m}\cdot\begin{cases}0,\quad n=2\text{ or }n>8,\\ -i\frac{\pi^{2}}{2}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{ijkl}\bigl{(}\delta_{ab}\delta_{cl}+\delta_{al}\delta_{bc}-\delta_{ac}\delta_{bl}\bigr{)},\quad n=4,\\ \frac{\pi^{3}}{4}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{abcijk},\quad n=6.\end{cases}$ (4.15)
Proof 4.11.
@@
$\mathscr{T}^{\chi}_{D}(u,v,w)=\frac{i\nu_{n-1}}{4}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}{}^{A}T_{ijk}\,\mathrm{Tr}\left(\gamma\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{i}\gamma^{j}\gamma^{k}\right).$ (4.16)
@@
$\begin{split}\mathscr{T}^{\chi}_{D}(u,v,w)&=-i\frac{\nu_{3}}{4}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{ijkl}\,\mathrm{Tr}(\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{l})\\ &=-\frac{i\nu_{3}}{4}\cdot 2^{m}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{ijkl}\bigl{(}\delta_{ab}\delta_{cl}+\delta_{al}\delta_{bc}-\delta_{ac}\delta_{bl}\bigr{)}.\end{split}$ (4.17)
@@
$\mathscr{T}^{\chi}_{D}(u,v,w)=\frac{\nu_{5}}{4}\cdot 2^{m}\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{abcijk}.$ (4.18)
@@
Proof 4.13.
First, we write the operator $\gamma\hat{u}\{D_{0},\hat{w}\}D_{0}|D_{0}|^{-n}$ as $(O_{1}+O_{2})L^{-m}$, where
$O_{1}=\gamma\hat{u}D_{0}\hat{w}D_{0},\quad O_{2}=\gamma\hat{u}\hat{w}D_{0}^{2}.$ (4.19)
@@
$\begin{split}F^{jk}=&u_{a}w_{b}\gamma\gamma^{a}\gamma^{(j|}\gamma^{b}\gamma^{|k)}=u_{a}(w_{k}\gamma\gamma^{a}\gamma^{j}+w_{j}\gamma\gamma^{a}\gamma^{k}-w_{b}\gamma\gamma^{a}\gamma^{b}\delta^{jk}),\\ G^{k}=&-u_{a}w_{bj}\gamma\gamma^{a}\gamma^{j}\gamma^{b}\gamma^{k},\\ H=&-\frac{1}{8}u_{a}w_{b}\gamma\gamma^{a}\gamma^{j}\gamma^{b}\gamma^{k}\gamma^{p}\gamma^{s}R_{jkps}=\frac{1}{4}u_{a}w_{b}\gamma\gamma^{a}\gamma^{j}\gamma^{b}\gamma^{d}\mathrm{Ric}_{jd}\\ =&\frac{1}{2}u_{a}w_{b}\mathrm{Ric}_{jb}\gamma\gamma^{a}\gamma^{j}-\frac{1}{4}u_{a}w_{b}\mathrm{Ric}_{jd}\gamma\gamma^{a}\gamma^{[j}\gamma^{d]}\gamma^{b}\\ =&\frac{1}{2}u_{a}w_{b}\mathrm{Ric}_{jb}\gamma\gamma^{a}\gamma^{j}-\frac{1}{4}u_{a}w_{b}R\gamma\gamma^{a}\gamma^{b}.\end{split}$ (4.20)
@@
$\begin{split}\mathscr{G}_{D_{0}}^{\chi}(u,w)&=\frac{\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(24H-12F^{ab}P_{ab}+4F^{ab}\mathrm{Ric}_{ab})\\ &=\frac{\nu_{n-1}}{6}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(6H-F^{ab}\mathrm{Ric}_{ab}).\end{split}$ (4.21)
Since $\mathrm{Tr}(\gamma\gamma^{a}\gamma^{b})=2i\delta_{n,2}\epsilon^{ab}$, from the form of the $F^{jk}$ and $H$ operators, we infer that $\mathscr{G}_{D_{0}}^{\chi}(u,w)$ vanishes identically for $n\neq 2$. For $n=2$ we are, strictly speaking, not allowed to use Corollary 2.3 directly, but we refer instead to Proposition 2.1 applied to $O=O_{1}$ and Theorem 2.2 applied to $E=O_{2}L^{-1}$. From the latter, we get $\mathscr{W}\!\mathrm{res}(EL^{-m+1})=0$, while the former, due to the fact that in $n=2$ dimensions we have $F^{jj}=0$, again leads to
$\begin{split}\mathscr{G}_{D_{0}}^{\chi}(u,w)&=\frac{\nu_{n-1}}{6}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(6H-F^{ab}\mathrm{Ric}_{ab}),\end{split}$ (4.22)
@@
$\begin{split}\mathscr{G}_{D_{0}}^{\chi}(u,w)&=i\frac{\nu_{1}}{3}\int_{M}\mathrm{dvol}_{g}\bigl{(}3u_{a}w_{b}\mathrm{Ric}_{jb}\epsilon^{aj}-\frac{3}{2}u_{a}w_{b}R\epsilon^{ab}-u_{a}w_{k}\epsilon^{aj}\mathrm{Ric}_{jk}\\ &\hskip 71.13188pt-u_{a}w_{j}\epsilon^{ak}\mathrm{Ric}_{jk}+u_{a}w_{b}\epsilon^{ab}\delta^{jk}\mathrm{Ric}_{jk}\bigr{)}\\ &=i\frac{\nu_{1}}{6}\int_{M}\mathrm{dvol}_{g}\bigl{[}\mathrm{2Ric}_{jb}\epsilon^{aj}-\epsilon^{ab}R\bigr{]}=0,\end{split}$ (4.23)
since for two-dimensional manifolds it is known that $\mathrm{Ric}=\frac{R}{2}g$.
@@
$\mathscr{G}_{D}^{\chi}(u,w)=\begin{cases}2\pi^{2}\int_{M}\mathrm{dvol}_{g}u_{a}\bigl{[}T_{ijk}^{0}(\varepsilon_{ijka}w_{bb}-\varepsilon_{ijkc}w_{ac}+\varepsilon_{ijkb}w_{ba})+w_{b}\varepsilon_{ijka}T_{ijk}^{c}\\ \hskip 71.13188pt+\frac{3}{4}w_{b}\varepsilon_{ijkl}T^{0}_{ijk}T^{0}_{abl}\bigr{]},\qquad n=4,\\ i\pi^{3}\int_{M}\mathrm{dvol}_{g}u_{a}\bigl{(}2w_{bc}T^{0}_{ijk}\varepsilon_{ijkabc}-w_{b}T^{c}_{ijk}\varepsilon_{ijkabc}\bigr{)},\qquad n=6,\\ 0,\qquad\text{otherwise}.\end{cases}$ (4.24)
Proof 4.15.
@@
For $n=4$, we have $\gamma B=-B\gamma=-\frac{i}{8}\varepsilon_{abcd}T_{abc}\gamma^{d}$, and the claim follows by a straightforward computation using Proposition 4.2. For $n=6$ we have $B\gamma=-\gamma B=-\frac{1}{48}T_{abc}\varepsilon_{abcijk}\gamma^{i}\gamma^{j}\gamma^{k}$and $\{\gamma^{d},\varepsilon_{abcijk}\gamma^{i}\gamma^{j}\gamma^{k}\}=6\varepsilon_{abcdjk}\gamma^{j}\gamma^{k}$, so that
$\mathrm{Tr}(\gamma^{d}\gamma^{e}\gamma^{f}\varepsilon_{abcijk}\gamma^{i}\gamma^{j}\gamma^{k})=-6\varepsilon_{abcdef}.$ (4.25)
This, together with Proposition 4.2 concludes the proof.
@@
$\mathscr{R}_{D}^{\chi}(f)=-\frac{{2^{m}}\nu_{3}}{8}\delta_{n,4}\int_{M}\mathrm{dvol}_{g}\varepsilon^{abcd}\partial_{a}T_{bcd},$ (4.26)
@@
Proof 4.17.
Since in this case we have $B=-\frac{i}{8}T_{abc}\gamma^{a}\gamma^{b}\gamma^{c}$ and $\chi=\gamma$ (the product of all gamma matrices), we have
$\begin{split}-\frac{i}{2}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(\chi\gamma^{a}B_{a})&=-\frac{i}{2}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\biggl{[}\gamma\gamma^{a}\biggl{(}-\frac{i}{8}\partial_{a}T_{bcd}\gamma^{b}\gamma^{c}\gamma^{d}\biggr{)}\biggr{]}\\ &=-\frac{(n-2)\nu_{n-1}}{16}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(\gamma\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d})\partial_{a}T_{bcd}\\ &=-{2^{m}}\frac{\nu_{n-1}(n-2)}{16}\delta_{n,4}\int_{M}\mathrm{dvol}_{g}\varepsilon^{abcd}\partial_{a}T_{bcd}\\ &=-{2^{m}}\frac{\nu_{3}}{8}\delta_{n,4}\int_{M}\varepsilon^{abcd}\partial_{a}T_{bcd},\end{split}$ (4.27)
@@
$\mathscr{W}\!\mathrm{res}(\chi D_{0}^{-n+2})=\frac{\nu_{n-1}}{24}(n-2)\int_{M}\mathrm{Tr}\bigl{[}\gamma(-3R-2R+4R)\bigr{]}=0.$ (4.28)
@@
$\mathscr{W}\!\mathrm{res}(u_{a}\gamma\gamma^{a}DD^{-2m})={2^{m}}\frac{i\nu_{3}}{4}\delta_{n,4}u_{a}T_{ijk}\varepsilon_{aijk}.$ (4.29)
4.2 The Hodge-Dirac operator
The spectral triple of the Hodge-Dirac operator is equipped with two distinct gradings. Before proceeding, we introduce a convenient notation:
$\gamma^{p}=-i(\lambda_{+}^{p}-\lambda_{-}^{p}),\qquad\qquad\tilde{\gamma}^{p}=\lambda^{p}_{+}+\lambda^{p}_{-}.$ (4.30)
Note that these operators are hermitian and satisify,
$\{\gamma^{p},\gamma^{s}\}=\{\tilde{\gamma}^{p},\tilde{\gamma}^{s}\}=2\delta^{ps},\qquad\{\tilde{\gamma}^{p},\gamma^{s}\}=0.$ (4.31)

### v2

@@ 4 Chiral spectral functionals
$\mathscr{W}\mathrm{res}\bigl(\chi D^{-n+2}\bigr)-\mathscr{W}\mathrm{res}\bigl(\chi D_{0}^{-n+2}\bigr)$
$\qquad=\frac{\nu_{n-1}}{24}(n-2)\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\bigl[\chi\bigl(-12B_{0}^{2}-12{\rm i}\gamma^{a}B_{a}+6{\rm i}\{\gamma^{a},B_{a}\}+3\{\gamma^{a},B\}\{\gamma^{a},B\}\bigr)\bigr]$
$\qquad=\frac{\nu_{n-1}}{8}(n-2)\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\bigl[\chi\bigl(-4B_{0}^{2}-4{\rm i}\gamma^{a}B_{a}+\{\gamma^{a},B\}\{\gamma^{a},B\}\bigr)\bigr]$
$\qquad=-\frac{\rm i}{2}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(\chi\gamma^{a}B_{a}).$
@@ 4.1 The spin Dirac operator
$\mathscr{g}_{D_{0}}^{\chi}(u,w)=\begin{cases}\displaystyle 4\pi{\rm i}{2^{m}}\int_{M}\mathrm{dvol}_{g}u_{a}w_{b}\varepsilon^{ab},&n=2,\\ 0,&n>2.\end{cases}$
Proof.
It follows from Proposition 2.1. ∎
@@
$\mathscr{T}_{D}^{\chi}(u,v,w)=2^{m}\cdot\begin{cases}0,&n=2\text{ or }n\geq 8,\\ \displaystyle-{\rm i}\frac{\pi^{2}}{2}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}&\\ \qquad{}\times{}^{A}T_{ijk}\varepsilon_{ijkl}\bigl(\delta_{ab}\delta_{cl}+\delta_{al}\delta_{bc}-\delta_{ac}\delta_{bl}\bigr),&n=4,\\ \displaystyle\frac{\pi^{3}}{4}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{abcijk},&n=6.\end{cases}$
Proof.
@@
$\mathscr{T}^{\chi}_{D}(u,v,w)=\frac{{\rm i}\nu_{n-1}}{4}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}{}^{A}T_{ijk}\mathrm{Tr}\bigl(\gamma\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{i}\gamma^{j}\gamma^{k}\bigr).$
@@
$\mathscr{T}^{\chi}_{D}(u,v,w)$ $=-{\rm i}\frac{\nu_{3}}{4}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{ijkl}\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{l}\bigr)$
$=-\frac{{\rm i}\nu_{3}}{4}\cdot 2^{m}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{ijkl}(\delta_{ab}\delta_{cl}+\delta_{al}\delta_{bc}-\delta_{ac}\delta_{bl}).$
@@
$\mathscr{T}^{\chi}_{D}(u,v,w)=\frac{\nu_{5}}{4}\cdot 2^{m}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}{}^{A}T_{ijk}\varepsilon_{abcijk}.$
@@
Proof.
[…]
$O_{1}=\gamma\hat{u}D_{0}\hat{w}D_{0},\qquad O_{2}=\gamma\hat{u}\hat{w}D_{0}^{2}.$
@@
$F^{jk}=u_{a}w_{b}\gamma\gamma^{a}\gamma^{(j|}\gamma^{b}\gamma^{|k)}=u_{a}\bigl(w_{k}\gamma\gamma^{a}\gamma^{j}+w_{j}\gamma\gamma^{a}\gamma^{k}-w_{b}\gamma\gamma^{a}\gamma^{b}\delta^{jk}\bigr),$
$G^{k}=-u_{a}w_{bj}\gamma\gamma^{a}\gamma^{j}\gamma^{b}\gamma^{k},$
$H=-\frac{1}{8}u_{a}w_{b}\gamma\gamma^{a}\gamma^{j}\gamma^{b}\gamma^{k}\gamma^{p}\gamma^{s}R_{jkps}=\frac{1}{4}u_{a}w_{b}\gamma\gamma^{a}\gamma^{j}\gamma^{b}\gamma^{d}\mathrm{Ric}_{jd}$
$\phantom{H}{}=\frac{1}{2}u_{a}w_{b}\mathrm{Ric}_{jb}\gamma\gamma^{a}\gamma^{j}-\frac{1}{4}u_{a}w_{b}\mathrm{Ric}_{jd}\gamma\gamma^{a}\gamma^{[j}\gamma^{d]}\gamma^{b}=\frac{1}{2}u_{a}w_{b}\mathrm{Ric}_{jb}\gamma\gamma^{a}\gamma^{j}-\frac{1}{4}u_{a}w_{b}R\gamma\gamma^{a}\gamma^{b}.$
@@
$\mathscr{G}_{D_{0}}^{\chi}(u,w)$ $=\frac{\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\bigl(24H-12F^{ab}P_{ab}+4F^{ab}\mathrm{Ric}_{ab}\bigr)$
$=\frac{\nu_{n-1}}{6}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\bigl(6H-F^{ab}\mathrm{Ric}_{ab}\bigr).$
Since $\mathrm{Tr}\bigl(\gamma\gamma^{a}\gamma^{b}\bigr)=2i\delta_{n,2}\epsilon^{ab}$, from the form of the $F^{jk}$ and $H$ operators, we infer that $\mathscr{G}_{D_{0}}^{\chi}(u,w)$ vanishes identically for $n\neq 2$. For $n=2$, we are, strictly speaking, not allowed to use Corollary 2.3 directly, but we refer instead to Proposition 2.1 applied to $O=O_{1}$ and Theorem 2.2 applied to $E=O_{2}L^{-1}$. From the latter, we get $\mathscr{W}\mathrm{res}\bigl(EL^{-m+1}\bigr)=0$, while the former, due to the fact that in $n=2$ dimensions we have $F^{jj}=0$, again leads to
$\mathscr{G}_{D_{0}}^{\chi}(u,w)=\frac{\nu_{n-1}}{6}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\bigl(6H-F^{ab}\mathrm{Ric}_{ab}\bigr),$
@@
$\mathscr{G}_{D_{0}}^{\chi}(u,w)={}$ ${\rm i}\frac{\nu_{1}}{3}\int_{M}\mathrm{dvol}_{g}\biggl(3u_{a}w_{b}\mathrm{Ric}_{jb}\epsilon^{aj}-\frac{3}{2}u_{a}w_{b}R\epsilon^{ab}-u_{a}w_{k}\epsilon^{aj}\mathrm{Ric}_{jk}$
$-u_{a}w_{j}\epsilon^{ak}\mathrm{Ric}_{jk}+u_{a}w_{b}\epsilon^{ab}\delta^{jk}\mathrm{Ric}_{jk}\biggr)$
$={}$ ${\rm i}\frac{\nu_{1}}{6}\int_{M}\mathrm{dvol}_{g}\bigl[\mathrm{2Ric}_{jb}\epsilon^{aj}-\epsilon^{ab}R\bigr]=0,$
since for two-dimensional manifolds it is known that $\mathrm{Ric}=\frac{R}{2}g$. ∎
@@
$\mathscr{G}_{D}^{\chi}(u,w)=\begin{cases}\displaystyle 2\pi^{2}\int_{M}\mathrm{dvol}_{g}u_{a}\biggl[T_{ijk}^{0}(\varepsilon_{ijka}w_{bb}-\varepsilon_{ijkc}w_{ac}+\varepsilon_{ijkb}w_{ba})&\\ \displaystyle\qquad{}+w_{b}\varepsilon_{ijka}T_{ijk}^{c}+\frac{3}{4}w_{b}\varepsilon_{ijkl}T^{0}_{ijk}T^{0}_{abl}\biggr],&n=4,\\ \displaystyle{\rm i}\pi^{3}\int_{M}\mathrm{dvol}_{g}u_{a}\bigl(2w_{bc}T^{0}_{ijk}\varepsilon_{ijkabc}-w_{b}T^{c}_{ijk}\varepsilon_{ijkabc}\bigr),&n=6,\\ 0,&\text{otherwise}.\end{cases}$
Proof.
@@
For $n=4$, we have $\gamma B=-B\gamma=-\frac{\rm i}{8}\varepsilon_{abcd}T_{abc}\gamma^{d}$, and the claim follows by a straightforward computation using Proposition 4.2. For $n=6$ we have $B\gamma=-\gamma B=-\frac{1}{48}T_{abc}\varepsilon_{abcijk}\gamma^{i}\gamma^{j}\gamma^{k}$ and $\bigl\{\gamma^{d},\varepsilon_{abcijk}\gamma^{i}\gamma^{j}\gamma^{k}\bigr\}=6\varepsilon_{abcdjk}\gamma^{j}\gamma^{k}$, so that $\mathrm{Tr}\bigl(\gamma^{d}\gamma^{e}\gamma^{f}\varepsilon_{abcijk}\gamma^{i}\gamma^{j}\gamma^{k}\bigr)=-6\varepsilon_{abcdef}$. This, together with Proposition 4.2 concludes the proof. ∎
@@
$\mathscr{R}_{D}^{\chi}(f)=-\frac{{2^{m}}\nu_{3}}{8}\delta_{n,4}\int_{M}\mathrm{dvol}_{g}\varepsilon^{abcd}\partial_{a}T_{bcd},$
@@
Proof.
Since in this case we have $B=-\frac{\rm i}{8}T_{abc}\gamma^{a}\gamma^{b}\gamma^{c}$ and $\chi=\gamma$ (the product of all gamma matrices), we have
$-\frac{\rm i}{2}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(\chi\gamma^{a}B_{a})$
$\qquad=-\frac{\rm i}{2}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\biggl[\gamma\gamma^{a}\biggl(-\frac{\rm i}{8}\partial_{a}T_{bcd}\gamma^{b}\gamma^{c}\gamma^{d}\biggr)\biggr]$
$\qquad=-\frac{(n-2)\nu_{n-1}}{16}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\bigl(\gamma\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}\bigr)\partial_{a}T_{bcd}$
$\qquad=-{2^{m}}\frac{\nu_{n-1}(n-2)}{16}\delta_{n,4}\int_{M}\mathrm{dvol}_{g}\varepsilon^{abcd}\partial_{a}T_{bcd}=-{2^{m}}\frac{\nu_{3}}{8}\delta_{n,4}\int_{M}\varepsilon^{abcd}\partial_{a}T_{bcd},$
@@
$\mathscr{W}\mathrm{res}\bigl(\chi D_{0}^{-n+2}\bigr)=\frac{\nu_{n-1}}{24}(n-2)\int_{M}\mathrm{Tr}\bigl[\gamma(-3R-2R+4R)\bigr]=0.$
@@
$\mathscr{W}\mathrm{res}\bigl(u_{a}\gamma\gamma^{a}DD^{-2m}\bigr)={2^{m}}\frac{{\rm i}\nu_{3}}{4}\delta_{n,4}u_{a}T_{ijk}\varepsilon_{aijk}.$
4.2 The Hodge–Dirac operator
The spectral triple of the Hodge–Dirac operator is equipped with two distinct gradings. Before proceeding, we introduce a convenient notation $\gamma^{p}=-{\rm i}(\lambda_{+}^{p}-\lambda_{-}^{p})$, $\tilde{\gamma}^{p}=\lambda^{p}_{+}+\lambda^{p}_{-}$. Note that these operators are hermitian and satisfy,
$\{\gamma^{p},\gamma^{s}\}=\{\tilde{\gamma}^{p},\tilde{\gamma}^{s}\}=2\delta^{ps},\qquad\{\tilde{\gamma}^{p},\gamma^{s}\}=0.$